OpenStudy (anonymous):

What is the slope of the line given by the equation below? y - 3 = 9(x + 10)

OpenStudy (anonymous):

y - 3 = 9x + 90--> y = 9x + 93. This is the standard form equation y = ax + b., where a is the slope.